Dear Collaborators, The official physics run for Cu+Cu will start tonight. Although the start-up has been quite bumpy, the machine made a good progress in the last week. The machine people believe that the machine now can deliver the beam routinely at minimum or higher luminosity planned/promised. At the planning meeting tomorrow, there will be some discussion how long the Cu+Cu run will last with this modified start-up schedule. Machine people will utilized some of the day shifts in the "physics running" period for machine development for higher luminosity and less background. The latest ZDC rates we had was 5-6K Hz , which is about ~4 times lower than PHENIX and STAR due to beta*. If you've been reading our e-log, you know more or less what we had to set-up and fix to make the experiment going. There are still some (mainly minor) issues we have to fix, but the experiment is quite ready for physics. The AC at FEH failed this afternoon. Hiro and Bob are still working on recovering the system form shutdown caused by the failure. When everything is recovered, we will start our first physics night with FS at 8 deg and MRS at 90 deg. Cheers, JH _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Tue Jan 11 20:45:28 2005
This archive was generated by hypermail 2.1.8 : Tue Jan 11 2005 - 20:45:52 EST